# Quasinormal spectra of a wormhole family: overtone features and a parameter-controlled redshift

### v1

@@ Appendix A QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
0 0 $0.83639-0.39420i$ $0.83639-0.39418i^{*}$- -- -- --
1 $1.78264-1.42784i$- -- -- -- --
1 0 $1.68973-0.28405i$ $1.68974-0.28404i^{*}$- -- -- --
1 $2.28101-1.18068i$- -- -- -- --
2 0 $2.64867-0.24369i$ $2.64856-0.24379i$ $0.30535863$ $1.86212-0.27064i$ $1.86214-0.27064i^{*}$ $0.24995736$
1 $3.09550-0.97920i$- -- $1.97161-0.71811i$- --
2 $3.73267-2.07179i$ $3.72699-2.07264i$ $0.22283351$ $2.55089-1.37091i$- --
3 0 $3.60155-3.36888i$ $3.62865-0.22143i$ $0.24259909$ $3.14398-0.31728i$ $3.14405-0.31724i$ $0.24589997$
1 $4.01130-0.86237i$- -- $3.18751-0.89811i$- --
2 $4.52123-1.81580i$ $4.52041-1.81625i$ $0.18309362$ $3.48149-1.49460i$ $3.48082-1.49494i$ $0.1574885$
4 0 $4.6164-0.20649i$ $4.61663-0.20641i$ $0.20468468$ $4.28394-0.28792i$ $4.28414-0.28787i$ $0.21497073$
1 $4.96092-0.78763i$ $4.95844-0.78954i$ $0.00019476$ $4.40055-0.92088i$- --
2 $5.40832-1.63672i$ $5.41382-1.63803i$ $0.15290777$ $4.59293-1.60512i$ $4.59135-1.60323i$ $0.1200131$
3 $5.94108-2.65940i$- -- $4.98385-2.34634i$- --
4 -- -- -- $5.60020-3.24404i$ $5.57299-3.21739i$ $0.12091509$
5 0 $5.60815-0.19523i$ $5.60859-0.19513i$ $0.1788866$ $5.35662-0.25860i$ $5.35702-0.25851i$ $0.18764768$
1 $5.92623-0.73455i$ $5.92481-0.73471i$ $0.00011546$ $5.53163-0.86711i$ $5.54394-0.87820i$ $1.795\times 10^{-5}$
2 $6.33667-1.50966i$ $6.33719-1.51452i$ $0.13574164$ $5.74869-1.60115i$ $5.74805-1.60159i$ $0.11547397$
3 $6.81719-2.45107i$- -- $5.91553-2.26906i$- --
4 -- -- -- $6.46920-3.21880i$ $6.41666-3.21904i$ $0.0932583$
6 0 $6.60210-0.18623i$ $6.60282-0.18610i$ $0.15997014$ $6.39998-0.23613i$ $6.40069-0.23600i$ $0.1668965$
1 $6.90042-0.69409i$ $6.90291-0.69404i$ $1.259\times 10^{-5}$ $6.60250-0.80713i$ $6.60367-0.81068i$ $2.414\times 10^{-5}$
2 $7.28536-1.41469i$ $7.28540-1.41346i$ $0.11879929$ $6.85764-1.53123i$ $6.85660-1.52978i$ $0.10977835$
3 $7.72908-2.28317i$- -- $7.21059-2.33977i$- --
4 $8.24197-3.28172i$ $8.25637-3.28542i$ $0.09481262$ $7.48045-3.19207i$ $7.50483-3.1369i$ $0.07558219$
7 0 $7.59739-0.17875i$ $7.59853-0.17859i$ $0.1454134$ $7.42853-0.21893i$ $7.42962-0.21877i$ $0.15090914$
1 $7.88016-0.66170i$ $7.88124-0.66145i$ $4.624\times 10^{-5}$ $7.64284-0.75603i$ $7.64344-0.75595i$ $6.13\times 10^{-6}$
2 $8.24579-1.34029i$ $8.24539-1.33931i$ $0.10773933$ $7.91941-1.45107i$ $7.91918-1.45030i$ $0.10293865$
3 $8.65746-2.16049i$- -- $8.22111-2.20332i$- --
4 $9.13470-3.08560i$ $9.10375-3.06571i$ $0.0848906$ $8.54449-3.11196i$ $8.54457-3.06192i$ $0.06786849$
@@
Table 9: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=4$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
@@
0 0 $0.86815-0.34320i$ $0.86815-0.34319i^{*}$- -- -- --
1 $2.04851-1.11401i$- -- -- -- --
2 $3.38964-1.91968i$ $3.39861-1.92833i^{*}$- -- -- --
1 0 $1.71264-0.22384i$ $1.71265-0.22383i^{*}$- -- -- --
1 $2.49623-0.91263i$- -- -- -- --
2 $3.64769-1.77155i$ $3.64832-1.77127i^{*}$- -- -- --
2 0 $2.66518-0.17757i$ $2.66523-0.17779i$ $0.29315965$ $1.89443-0.24761i$ $1.89445-0.24761i^{*}$ $0.2266797$
1 $3.24903-0.73077i$- -- $2.04700-0.61385i$- --
2 $4.15868-1.54933i$ $4.15833-1.54843i$ $0.20476605$ $2.78063-1.07127i$ $2.77464-1.07131i^{*}$ $0.30966731$
3 $5.28869-2.41854i$- -- -- -- --
4 $6.52829-3.26724i$ $6.53412-3.27928i$ $0.40308849$- -- --
3 0 $3.64016-0.15237i$ $3.64024-0.15233i$ $0.22714669$ $3.18363-0.28391i$ $3.18372-0.28388i$ $0.23197406$
1 $4.12259-0.61681i$ $4.14739-0.63000i$ $0.00015229$ $3.25390-0.76629i$- --
2 $4.86332-1.34256i$ $4.86269-1.34213i$ $0.18258948$ $3.65647-1.17197i$ $3.65640-1.17160i$ $0.16536179$
3 $5.81271-2.19639i$- -- $4.52375-1.72969i$- --
4 $6.94662-3.06045i$ $6.95026-3.06342i$ $0.12904273$ $5.60501-2.47686i$ $5.60406-2.47709i$ $0.15103107$
4 0 $4.62419-0.13598i$ $4.62440-0.13592i$ $0.18790256$ $4.32457-0.23666i$ $4.32477-0.23661i$ $0.20448213$
1 $5.04395-0.54241i$ $5.04408-0.54139i$ $1.859\times 10^{-5}$ $4.47738-0.78946i$- --
2 $5.67924-1.18324i$ $5.67895-1.18265i$ $0.15644501$ $4.71713-1.30051i$ $4.71834-1.29988i$ $0.10481977$
3 $6.51135-1.97082i$- -- $5.30511-1.77828i$- --
4 $7.49665-2.83332i$ $7.49472-2.83864i$ $0.11727697$ $6.23010-2.39260i$ $6.22995-2.38931i$ $0.13343803$
5 0 $5.61290-0.12418i$ $5.61332-0.12410i$ $0.16150451$ $5.39035-0.19645i$ $5.39075-0.19637i$ $0.17505116$
1 $5.98915-0.48994i$ $5.98930-0.48917i$ $7.8\times 10^{-6}$ $5.62591-0.70862i$ $5.62610-0.71015i$ $4.315\times 10^{-5}$
2 $6.55411-1.06444i$ $6.55411-1.06349i$ $0.13585498$ $5.87683-1.32855i$ $5.87679-1.32565i$ $0.09890309$
3 $7.28065-1.79073i$- -- $6.26018-1.85524i$- --
4 $8.15775-2.61273i$ $8.15569-2.60335i$ $0.10450563$ $6.96511-2.38395i$ $6.9809-2.36109i$ $0.10370454$
6 0 $6.60441-0.11513i$ $6.60513-0.11501i$ $0.14239417$ $6.42671-0.16857i$ $6.42738-0.16847i$ $0.15284811$
1 $6.94825-0.45059i$ $6.94923-0.45040i$ $0.00018714$ $6.69168-0.62009i$ $6.69217-0.62012i$ $0.00011249$
2 $7.46242-0.97399i$ $7.46683-0.96882i$ $0.1139611$ $7.01628-1.24539i$ $7.01625-1.24492i$ $0.10544664$
3 $8.11603-1.64203i$- -- $7.35083-1.88535i$- --
4 $8.90307-2.41503i$ $8.79939-2.40900i$ $0.06811125$ $7.84927-2.43946i$ $7.84363-2.43251i$ $0.08016603$
7 0 $7.59771-0.10787i$ $7.59880-0.10775i$ $0.12789715$ $7.44938-0.14872i$ $7.45045-0.14859i$ $0.13601879$
1 $7.91626-0.41968i$ $7.91704-0.41916i$ $8.235\times 10^{-5}$ $7.71959-0.55112i$ $7.72040-0.55064i$ $3.859\times 10^{-5}$
2 $8.39152-0.90296i$ $8.39220-0.90173i$ $0.10758787$ $8.08261-1.12967i$ $8.08311-1.12881i$ $0.10269501$
3 $8.99190-1.52108i$ $8.93263-1.51227i$ $0.00374277$ $8.46623-1.80756i$- --
4 $9.70803-2.24495i$ $9.69532-2.23123i$ $0.09946345$ $8.87902-2.45825i$ $8.88319-2.49385i$ $0.06576324$
@@
Table 10: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=6$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
@@
0 0 $0.87997-0.32519i$ $0.87998-0.32518i^{*}$- -- -- --
1 $2.14466-0.97491i$- -- -- -- --
2 $3.57108-1.59477i$ $3.57054-1.59493i^{*}$- -- -- --
1 0 $1.72193-0.20105i$ $1.72194-0.20104i^{*}$- -- -- --
1 $2.57737-0.79070i$- -- -- -- --
2 $3.82365-1.47070i$ $3.82587-1.46941i^{*}$- -- -- --
2 0 $2.67187-0.15181i$ $2.67251-0.15238i$ $0.29083766$ $1.91003-0.23457i$ $1.91005-0.23456i^{*}$ $0.21776703$
1 $3.30697-0.61908i$- -- $2.08571-0.56896i$- --
2 $4.31495-1.28267i$ $4.31427-1.28190i$ $0.23210154$ $2.87446-0.93747i$ $2.87405-0.93741i^{*}$ $0.29184553$
3 $5.55107-1.96109i$- -- $4.04950-1.50405i$- --
4 $6.88381-2.61925i$ $6.88773-2.61918i$ $0.25385346$ $5.37019-2.11453i$ $5.36933-2.11477i^{*}$ $0.1670686$
3 0 $3.64456-0.12512i$ $3.64471-0.12510i$ $0.22202515$ $3.19594-0.26645i$ $3.19603-0.26648i$ $0.22253281$
1 $4.16276-0.50814i$ $4.17354-0.50844i$ $0.00037032$ $3.28608-0.69780i$- --
2 $4.98966-1.10188i$ $4.99007-1.10179i$ $0.18332356$ $3.73814-1.01798i$ $3.73866-1.01791i$ $0.17541431$
3 $6.06575-1.77885i$- -- $4.69073-1.44514i$- --
4 $7.29017-2.45688i$ $7.29374-2.46700i$ $0.13255104$ $5.86562-2.02004i$ $5.8667-2.02183i$ $0.15024292$
4 0 $4.62667-0.10807i$ $4.62688-0.10801i$ $0.18186996$ $4.33741-0.21472i$ $4.33761-0.21467i$ $0.19968836$
1 $5.07164-0.43530i$ $5.07146-0.43473i$ $5.0481\times 10^{-6}$ $4.49808-0.72204i$- --
2 $5.77715-0.95771i$ $5.77702-0.95720i$ $0.15714858$ $4.76792-1.13746i$ $4.76742-1.13632i$ $0.10174903$
3 $6.71074-1.59843i$- -- $5.44241-1.46518i$- --
4 $7.81676-2.27620i$ $7.81867-2.27681i$ $0.11845874$ $6.47562-1.94993i$ $6.47398-1.95101i$ $0.13691266$
5 0 $5.61380-0.09608i$ $5.61421-0.09601i$ $0.15498730$ $5.40063-0.17075i$ $5.40103-0.17067i$ $0.16992681$
1 $6.00767-0.38426i$ $6.00786-0.38394i$ $1.7893\times 10^{-5}$ $5.65481-0.64034i$- --
2 $6.62899-0.84805i$ $6.62886-0.84759i$ $0.13681384$ $5.90429-1.18229i$ $5.90559-1.17820i$ $0.08731672$
3 $7.45181-1.43958i$ $7.45249-1.39775i$ $0.00057971$ $6.34149-1.56146i$- --
4 $8.44520-2.09680i$ $8.44251-2.10081i$ $0.11003456$ $7.17592-1.93339i$ $7.18692-1.93325i$ $0.11707173$
6 0 $6.60399-0.08711i$ $6.60469-0.08702i$ $0.13560445$ $6.43410-0.14111i$ $6.43478-0.14102i$ $0.14713892$
1 $6.95973-0.34643i$ $6.96018-0.34614i$ $1.3489\times 10^{-5}$ $6.72107-0.54020i$ $6.72167-0.53992i$ $8.009\times 10^{-5}$
2 $7.51913-0.76397i$ $7.51916-0.76313i$ $0.12066340$ $7.05523-1.11725i$ $7.05465-1.11561i$ $0.09874688$
3 $8.25731-1.30704i$ $8.23586-1.33897i$ $0.00457058$ $7.39302-1.62858i$- --
4 $9.15431-1.93145i$ $9.15519-1.93202i$ $0.10141119$ $7.98533-1.99226i$ $7.98248-1.98655i$ $0.08462117$
7 0 $7.59621-0.08009i$ $7.59732-0.07999i$ $0.12093846$ $7.45434-0.12059i$ $7.45540-0.12047i$ $0.12992236$
1 $7.92220-0.31715i$ $7.92296-0.31677i$ $3.2848\times 10^{-5}$ $7.74342-0.46352i$ $7.74431-0.46321i$ $1.434\times 10^{-5}$
2 $8.43378-0.69798i$ $8.43393-0.69726i$ $0.10853332$ $8.13502-0.98974i$ $8.13560-0.98861i$ $0.10233494$
3 $9.10647-1.19796i$ $9.10056-1.19398i$ $0.00094986$ $8.51244-1.59506i$- --
4 $9.92416-1.78541i$ $9.92248-1.77980i$ $0.09058615$ $8.94988-2.06737i$ $8.95234-2.06370i$ $0.06334415$
@@
Table 11: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=8$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]

### v2

@@ Appendix B QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
0 0 0.836390.39420i0.83639-0.39420i $0.83639-0.39418i^{*}$ - - - -
1 1.782641.42784i1.78264-1.42784i - - - - -
1 0 1.689730.28405i1.68973-0.28405i $1.68974-0.28404i^{*}$ - - - -
1 2.281011.18068i2.28101-1.18068i - - - - -
2 0 2.648670.24369i2.64867-0.24369i 2.648560.24379i2.64856-0.24379i $0.30535863$ 1.862120.27064i1.86212-0.27064i $1.86214-0.27064i^{*}$ $0.24995736$
1 3.095500.97920i3.09550-0.97920i - - 1.971610.71811i1.97161-0.71811i - -
2 3.732672.07179i3.73267-2.07179i 3.726992.07264i3.72699-2.07264i $0.22283351$ 2.550891.37091i2.55089-1.37091i - -
3 0 3.601553.36888i3.60155-3.36888i 3.628650.22143i3.62865-0.22143i $0.24259909$ 3.143980.31728i3.14398-0.31728i 3.144050.31724i3.14405-0.31724i $0.24589997$
1 4.011300.86237i4.01130-0.86237i - - 3.187510.89811i3.18751-0.89811i - -
2 4.521231.81580i4.52123-1.81580i 4.520411.81625i4.52041-1.81625i $0.18309362$ 3.481491.49460i3.48149-1.49460i 3.480821.49494i3.48082-1.49494i $0.1574885$
4 0 4.61640.20649i4.6164-0.20649i 4.616630.20641i4.61663-0.20641i $0.20468468$ 4.283940.28792i4.28394-0.28792i 4.284140.28787i4.28414-0.28787i $0.21497073$
1 4.960920.78763i4.96092-0.78763i 4.958440.78954i4.95844-0.78954i $0.00019476$ 4.400550.92088i4.40055-0.92088i - -
2 5.408321.63672i5.40832-1.63672i 5.413821.63803i5.41382-1.63803i $0.15290777$ 4.592931.60512i4.59293-1.60512i 4.591351.60323i4.59135-1.60323i $0.1200131$
3 5.941082.65940i5.94108-2.65940i - - 4.983852.34634i4.98385-2.34634i - -
4 - - - 5.600203.24404i5.60020-3.24404i 5.572993.21739i5.57299-3.21739i $0.12091509$
5 0 5.608150.19523i5.60815-0.19523i 5.608590.19513i5.60859-0.19513i $0.1788866$ 5.356620.25860i5.35662-0.25860i 5.357020.25851i5.35702-0.25851i $0.18764768$
1 5.926230.73455i5.92623-0.73455i 5.924810.73471i5.92481-0.73471i $0.00011546$ 5.531630.86711i5.53163-0.86711i 5.543940.87820i5.54394-0.87820i $1.795\times 10^{-5}$
2 6.336671.50966i6.33667-1.50966i 6.337191.51452i6.33719-1.51452i $0.13574164$ 5.748691.60115i5.74869-1.60115i 5.748051.60159i5.74805-1.60159i $0.11547397$
3 6.817192.45107i6.81719-2.45107i - - 5.915532.26906i5.91553-2.26906i - -
4 - - - 6.469203.21880i6.46920-3.21880i 6.416663.21904i6.41666-3.21904i $0.0932583$
6 0 6.602100.18623i6.60210-0.18623i 6.602820.18610i6.60282-0.18610i $0.15997014$ 6.399980.23613i6.39998-0.23613i 6.400690.23600i6.40069-0.23600i $0.1668965$
1 6.900420.69409i6.90042-0.69409i 6.902910.69404i6.90291-0.69404i $1.259\times 10^{-5}$ 6.602500.80713i6.60250-0.80713i 6.603670.81068i6.60367-0.81068i $2.414\times 10^{-5}$
2 7.285361.41469i7.28536-1.41469i 7.285401.41346i7.28540-1.41346i $0.11879929$ 6.857641.53123i6.85764-1.53123i 6.856601.52978i6.85660-1.52978i $0.10977835$
3 7.729082.28317i7.72908-2.28317i - - 7.210592.33977i7.21059-2.33977i - -
4 8.241973.28172i8.24197-3.28172i 8.256373.28542i8.25637-3.28542i $0.09481262$ 7.480453.19207i7.48045-3.19207i 7.504833.1369i7.50483-3.1369i $0.07558219$
7 0 7.597390.17875i7.59739-0.17875i 7.598530.17859i7.59853-0.17859i $0.1454134$ 7.428530.21893i7.42853-0.21893i 7.429620.21877i7.42962-0.21877i $0.15090914$
1 7.880160.66170i7.88016-0.66170i 7.881240.66145i7.88124-0.66145i $4.624\times 10^{-5}$ 7.642840.75603i7.64284-0.75603i 7.643440.75595i7.64344-0.75595i $6.13\times 10^{-6}$
2 8.245791.34029i8.24579-1.34029i 8.245391.33931i8.24539-1.33931i $0.10773933$ 7.919411.45107i7.91941-1.45107i 7.919181.45030i7.91918-1.45030i $0.10293865$
3 8.657462.16049i8.65746-2.16049i - - 8.221112.20332i8.22111-2.20332i - -
4 9.134703.08560i9.13470-3.08560i 9.103753.06571i9.10375-3.06571i $0.0848906$ 8.544493.11196i8.54449-3.11196i 8.544573.06192i8.54457-3.06192i $0.06786849$
@@
Table 9: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=4$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
@@
0 0 0.868150.34320i0.86815-0.34320i $0.86815-0.34319i^{*}$ - - - -
1 2.048511.11401i2.04851-1.11401i - - - - -
2 3.389641.91968i3.38964-1.91968i $3.39861-1.92833i^{*}$ - - - -
1 0 1.712640.22384i1.71264-0.22384i $1.71265-0.22383i^{*}$ - - - -
1 2.496230.91263i2.49623-0.91263i - - - - -
2 3.647691.77155i3.64769-1.77155i $3.64832-1.77127i^{*}$ - - - -
2 0 2.665180.17757i2.66518-0.17757i 2.665230.17779i2.66523-0.17779i $0.29315965$ 1.894430.24761i1.89443-0.24761i $1.89445-0.24761i^{*}$ $0.2266797$
1 3.249030.73077i3.24903-0.73077i - - 2.047000.61385i2.04700-0.61385i - -
2 4.158681.54933i4.15868-1.54933i 4.158331.54843i4.15833-1.54843i $0.20476605$ 2.780631.07127i2.78063-1.07127i $2.77464-1.07131i^{*}$ $0.30966731$
3 5.288692.41854i5.28869-2.41854i - - - - -
4 6.528293.26724i6.52829-3.26724i 6.534123.27928i6.53412-3.27928i $0.40308849$ - - -
3 0 3.640160.15237i3.64016-0.15237i 3.640240.15233i3.64024-0.15233i $0.22714669$ 3.183630.28391i3.18363-0.28391i 3.183720.28388i3.18372-0.28388i $0.23197406$
1 4.122590.61681i4.12259-0.61681i 4.147390.63000i4.14739-0.63000i $0.00015229$ 3.253900.76629i3.25390-0.76629i - -
2 4.863321.34256i4.86332-1.34256i 4.862691.34213i4.86269-1.34213i $0.18258948$ 3.656471.17197i3.65647-1.17197i 3.656401.17160i3.65640-1.17160i $0.16536179$
3 5.812712.19639i5.81271-2.19639i - - 4.523751.72969i4.52375-1.72969i - -
4 6.946623.06045i6.94662-3.06045i 6.950263.06342i6.95026-3.06342i $0.12904273$ 5.605012.47686i5.60501-2.47686i 5.604062.47709i5.60406-2.47709i $0.15103107$
4 0 4.624190.13598i4.62419-0.13598i 4.624400.13592i4.62440-0.13592i $0.18790256$ 4.324570.23666i4.32457-0.23666i 4.324770.23661i4.32477-0.23661i $0.20448213$
1 5.043950.54241i5.04395-0.54241i 5.044080.54139i5.04408-0.54139i $1.859\times 10^{-5}$ 4.477380.78946i4.47738-0.78946i - -
2 5.679241.18324i5.67924-1.18324i 5.678951.18265i5.67895-1.18265i $0.15644501$ 4.717131.30051i4.71713-1.30051i 4.718341.29988i4.71834-1.29988i $0.10481977$
3 6.511351.97082i6.51135-1.97082i - - 5.305111.77828i5.30511-1.77828i - -
4 7.496652.83332i7.49665-2.83332i 7.494722.83864i7.49472-2.83864i $0.11727697$ 6.230102.39260i6.23010-2.39260i 6.229952.38931i6.22995-2.38931i $0.13343803$
5 0 5.612900.12418i5.61290-0.12418i 5.613320.12410i5.61332-0.12410i $0.16150451$ 5.390350.19645i5.39035-0.19645i 5.390750.19637i5.39075-0.19637i $0.17505116$
1 5.989150.48994i5.98915-0.48994i 5.989300.48917i5.98930-0.48917i $7.8\times 10^{-6}$ 5.625910.70862i5.62591-0.70862i 5.626100.71015i5.62610-0.71015i $4.315\times 10^{-5}$
2 6.554111.06444i6.55411-1.06444i 6.554111.06349i6.55411-1.06349i $0.13585498$ 5.876831.32855i5.87683-1.32855i 5.876791.32565i5.87679-1.32565i $0.09890309$
3 7.280651.79073i7.28065-1.79073i - - 6.260181.85524i6.26018-1.85524i - -
4 8.157752.61273i8.15775-2.61273i 8.155692.60335i8.15569-2.60335i $0.10450563$ 6.965112.38395i6.96511-2.38395i 6.98092.36109i6.9809-2.36109i $0.10370454$
6 0 6.604410.11513i6.60441-0.11513i 6.605130.11501i6.60513-0.11501i $0.14239417$ 6.426710.16857i6.42671-0.16857i 6.427380.16847i6.42738-0.16847i $0.15284811$
1 6.948250.45059i6.94825-0.45059i 6.949230.45040i6.94923-0.45040i $0.00018714$ 6.691680.62009i6.69168-0.62009i 6.692170.62012i6.69217-0.62012i $0.00011249$
2 7.462420.97399i7.46242-0.97399i 7.466830.96882i7.46683-0.96882i $0.1139611$ 7.016281.24539i7.01628-1.24539i 7.016251.24492i7.01625-1.24492i $0.10544664$
3 8.116031.64203i8.11603-1.64203i - - 7.350831.88535i7.35083-1.88535i - -
4 8.903072.41503i8.90307-2.41503i 8.799392.40900i8.79939-2.40900i $0.06811125$ 7.849272.43946i7.84927-2.43946i 7.843632.43251i7.84363-2.43251i $0.08016603$
7 0 7.597710.10787i7.59771-0.10787i 7.598800.10775i7.59880-0.10775i $0.12789715$ 7.449380.14872i7.44938-0.14872i 7.450450.14859i7.45045-0.14859i $0.13601879$
1 7.916260.41968i7.91626-0.41968i 7.917040.41916i7.91704-0.41916i $8.235\times 10^{-5}$ 7.719590.55112i7.71959-0.55112i 7.720400.55064i7.72040-0.55064i $3.859\times 10^{-5}$
2 8.391520.90296i8.39152-0.90296i 8.392200.90173i8.39220-0.90173i $0.10758787$ 8.082611.12967i8.08261-1.12967i 8.083111.12881i8.08311-1.12881i $0.10269501$
3 8.991901.52108i8.99190-1.52108i 8.932631.51227i8.93263-1.51227i $0.00374277$ 8.466231.80756i8.46623-1.80756i - -
4 9.708032.24495i9.70803-2.24495i 9.695322.23123i9.69532-2.23123i $0.09946345$ 8.879022.45825i8.87902-2.45825i 8.883192.49385i8.88319-2.49385i $0.06576324$
@@
Table 10: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=6$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
@@
0 0 0.879970.32519i0.87997-0.32519i $0.87998-0.32518i^{*}$ - - - -
1 2.144660.97491i2.14466-0.97491i - - - - -
2 3.571081.59477i3.57108-1.59477i $3.57054-1.59493i^{*}$ - - - -
1 0 1.721930.20105i1.72193-0.20105i $1.72194-0.20104i^{*}$ - - - -
1 2.577370.79070i2.57737-0.79070i - - - - -
2 3.823651.47070i3.82365-1.47070i $3.82587-1.46941i^{*}$ - - - -
2 0 2.671870.15181i2.67187-0.15181i 2.672510.15238i2.67251-0.15238i $0.29083766$ 1.910030.23457i1.91003-0.23457i $1.91005-0.23456i^{*}$ $0.21776703$
1 3.306970.61908i3.30697-0.61908i - - 2.085710.56896i2.08571-0.56896i - -
2 4.314951.28267i4.31495-1.28267i 4.314271.28190i4.31427-1.28190i $0.23210154$ 2.874460.93747i2.87446-0.93747i $2.87405-0.93741i^{*}$ $0.29184553$
3 5.551071.96109i5.55107-1.96109i - - 4.049501.50405i4.04950-1.50405i - -
4 6.883812.61925i6.88381-2.61925i 6.887732.61918i6.88773-2.61918i $0.25385346$ 5.370192.11453i5.37019-2.11453i $5.36933-2.11477i^{*}$ $0.1670686$
3 0 3.644560.12512i3.64456-0.12512i 3.644710.12510i3.64471-0.12510i $0.22202515$ 3.195940.26645i3.19594-0.26645i 3.196030.26648i3.19603-0.26648i $0.22253281$
1 4.162760.50814i4.16276-0.50814i 4.173540.50844i4.17354-0.50844i $0.00037032$ 3.286080.69780i3.28608-0.69780i - -
2 4.989661.10188i4.98966-1.10188i 4.990071.10179i4.99007-1.10179i $0.18332356$ 3.738141.01798i3.73814-1.01798i 3.738661.01791i3.73866-1.01791i $0.17541431$
3 6.065751.77885i6.06575-1.77885i - - 4.690731.44514i4.69073-1.44514i - -
4 7.290172.45688i7.29017-2.45688i 7.293742.46700i7.29374-2.46700i $0.13255104$ 5.865622.02004i5.86562-2.02004i 5.86672.02183i5.8667-2.02183i $0.15024292$
4 0 4.626670.10807i4.62667-0.10807i 4.626880.10801i4.62688-0.10801i $0.18186996$ 4.337410.21472i4.33741-0.21472i 4.337610.21467i4.33761-0.21467i $0.19968836$
1 5.071640.43530i5.07164-0.43530i 5.071460.43473i5.07146-0.43473i $5.0481\times 10^{-6}$ 4.498080.72204i4.49808-0.72204i - -
2 5.777150.95771i5.77715-0.95771i 5.777020.95720i5.77702-0.95720i $0.15714858$ 4.767921.13746i4.76792-1.13746i 4.767421.13632i4.76742-1.13632i $0.10174903$
3 6.710741.59843i6.71074-1.59843i - - 5.442411.46518i5.44241-1.46518i - -
4 7.816762.27620i7.81676-2.27620i 7.818672.27681i7.81867-2.27681i $0.11845874$ 6.475621.94993i6.47562-1.94993i 6.473981.95101i6.47398-1.95101i $0.13691266$
5 0 5.613800.09608i5.61380-0.09608i 5.614210.09601i5.61421-0.09601i $0.15498730$ 5.400630.17075i5.40063-0.17075i 5.401030.17067i5.40103-0.17067i $0.16992681$
1 6.007670.38426i6.00767-0.38426i 6.007860.38394i6.00786-0.38394i $1.7893\times 10^{-5}$ 5.654810.64034i5.65481-0.64034i - -
2 6.628990.84805i6.62899-0.84805i 6.628860.84759i6.62886-0.84759i $0.13681384$ 5.904291.18229i5.90429-1.18229i 5.905591.17820i5.90559-1.17820i $0.08731672$
3 7.451811.43958i7.45181-1.43958i 7.452491.39775i7.45249-1.39775i $0.00057971$ 6.341491.56146i6.34149-1.56146i - -
4 8.445202.09680i8.44520-2.09680i 8.442512.10081i8.44251-2.10081i $0.11003456$ 7.175921.93339i7.17592-1.93339i 7.186921.93325i7.18692-1.93325i $0.11707173$
6 0 6.603990.08711i6.60399-0.08711i 6.604690.08702i6.60469-0.08702i $0.13560445$ 6.434100.14111i6.43410-0.14111i 6.434780.14102i6.43478-0.14102i $0.14713892$
1 6.959730.34643i6.95973-0.34643i 6.960180.34614i6.96018-0.34614i $1.3489\times 10^{-5}$ 6.721070.54020i6.72107-0.54020i 6.721670.53992i6.72167-0.53992i $8.009\times 10^{-5}$
2 7.519130.76397i7.51913-0.76397i 7.519160.76313i7.51916-0.76313i $0.12066340$ 7.055231.11725i7.05523-1.11725i 7.054651.11561i7.05465-1.11561i $0.09874688$
3 8.257311.30704i8.25731-1.30704i 8.235861.33897i8.23586-1.33897i $0.00457058$ 7.393021.62858i7.39302-1.62858i - -
4 9.154311.93145i9.15431-1.93145i 9.155191.93202i9.15519-1.93202i $0.10141119$ 7.985331.99226i7.98533-1.99226i 7.982481.98655i7.98248-1.98655i $0.08462117$
7 0 7.596210.08009i7.59621-0.08009i 7.597320.07999i7.59732-0.07999i $0.12093846$ 7.454340.12059i7.45434-0.12059i 7.455400.12047i7.45540-0.12047i $0.12992236$
1 7.922200.31715i7.92220-0.31715i 7.922960.31677i7.92296-0.31677i $3.2848\times 10^{-5}$ 7.743420.46352i7.74342-0.46352i 7.744310.46321i7.74431-0.46321i $1.434\times 10^{-5}$
2 8.433780.69798i8.43378-0.69798i 8.433930.69726i8.43393-0.69726i $0.10853332$ 8.135020.98974i8.13502-0.98974i 8.135600.98861i8.13560-0.98861i $0.10233494$
3 9.106471.19796i9.10647-1.19796i 9.100561.19398i9.10056-1.19398i $0.00094986$ 8.512441.59506i8.51244-1.59506i - -
4 9.924161.78541i9.92416-1.78541i 9.922481.77980i9.92248-1.77980i $0.09058615$ 8.949882.06737i8.94988-2.06737i 8.952342.06370i8.95234-2.06370i $0.06334415$
@@
Table 11: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=8$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]